OpenStudy (anonymous):

how many solutions are there to the following system of equations? 3x-9y=0 and -x+3y=-3

OpenStudy (freckles):

Well I would find the number by attempting to solve the system.

OpenStudy (freckles):

You easily solve 3x-9y=0 for x

OpenStudy (freckles):

then you can input that into the other equation -x+3y=-3 and try to solve for y

OpenStudy (anonymous):

still confused

OpenStudy (freckles):

can you solve 3x-9y=0 for x?

OpenStudy (anonymous):

no i am lost

OpenStudy (anonymous):

would you add nine to both sides?

OpenStudy (freckles):

3x-9y=0 Undo the subtraction of 9y by adding 9y (add on both sides)

OpenStudy (freckles):

3x-9y=0 +9y +9y ------------ 3x=9y you would add 9y just like i did right here

OpenStudy (freckles):

now can you get x by itself?

OpenStudy (freckles):

Hint: think dividing something on both sides

OpenStudy (anonymous):

divide both sides by 3?

OpenStudy (freckles):

x=9y/3 x=3y since 9/3=3

OpenStudy (anonymous):

x=3?

OpenStudy (freckles):

the other equation is -x+3y=-3 right?

OpenStudy (freckles):

so replace the x there with 3y

OpenStudy (freckles):

-(3y)+3y=-3 if you end up with the same thing on both sides you have infinitely many answers if you end up with something different on both sides you have 0 answers if you end up with something like y=a number you have one answer

OpenStudy (anonymous):

so it will be 0? am i right?

OpenStudy (freckles):

yes because -3y+3y=0 and 0 doesn't equal -3 right?

OpenStudy (freckles):

gj

OpenStudy (anonymous):

yeah!!!!!! thank you so much

OpenStudy (freckles):

one more thing: 3x-9y=0 -x+3y=-3 ------we can write both in y=mx+b form 3x-9y=0 add 9y on both sides 9y=3x divide by 9 on both sides y=3/9 x y=1/3 x ther other equation is -x+3y=-3 3y=x-3 I add x on both sides here y=1/3 x -3/3 I divided 3 on both sides y=1/3 x - 1 I simplified 3/3=1 now comparing y=1/3 x y=1/3 x - 1 They both have the same slope which is 1/3 So if two lines have the same slope, that means 1 of 2 things 1) same line if they have the same y-intercept 2) parallel if they have different y-intercept The case here is 2) since the first equation has y-intercept 0 and second equation has y-intercept -1 Parallel lines never intersect so that means they will not have s "solution" The solution is the set in which the lines intersect

OpenStudy (freckles):

That is just another way to look it. The possible solutions to a system of 2 linear equations with 2 variables is: 1) same slope and same y-intercept => infinitely many solutions (since same line) 2) same slope and different y-intercept=> 0 solutions (since parallel) 3) different slope=> 1 solution